OpenStudy (anonymous):

how do you do log(8^8 * 5^5)

OpenStudy (brunsonni):

When multiplying exponents you multiply the constants and add the exponents so that would be 40^13

OpenStudy (anonymous):

so it would be log(40^13)?

OpenStudy (anonymous):

@brunsonni You may want to check that with a calculator... That's not quite accurate

OpenStudy (anonymous):

wait so how do i do it?

OpenStudy (anonymous):

You are right that if the base was the same, you could add the exponents Ex: 2^4*2^2=2*2*2*2*2*2=2^6 But you can't just combine everything with different bases and powers :)

OpenStudy (anonymous):

As for this problem, the easiest way is to use a calculator. The mathier way is to use the properties of logarithms. In this case, you have to know that \[\log_{}(a*b )=\log(a)+\log(b)\] and that \[\log(a^b)=b \times \log(a)\]

OpenStudy (anonymous):

Can you figure it out from here?

OpenStudy (anonymous):

wait it says to expand the problem completely

OpenStudy (anonymous):

right, so you can use the properties I listed above to expand it. You'll have to use both and I recommend using them in the order I listed